OpenStudy (darkbluechocobo):

Classify the system and determine the number of solutions.Check all that apply. x + 4 = y 5y = 5x + 35 infinitely many solutions inconsistent no solution consistent I put inconsistent and no solution am I right?

OpenStudy (jdoe0001):

hehehe I saw that =P

OpenStudy (darkbluechocobo):

xD yes a facepalm moment

OpenStudy (jdoe0001):

I was about to say something about the $, sounded like an expensive matter or something

OpenStudy (darkbluechocobo):

xD all it is, is system of equations...which I hate ._.

OpenStudy (jdoe0001):

so... what makes you think is "inconsistent"?

OpenStudy (darkbluechocobo):

well... the 2nd equation after its all done would be y=x+6 and it doesnt match x+4=y

OpenStudy (jdoe0001):

well. do they have to match?

OpenStudy (jdoe0001):

the idea of a system of equations is that they're not supposed to match

OpenStudy (darkbluechocobo):

I guess not

OpenStudy (darkbluechocobo):

but it would be nice

OpenStudy (jdoe0001):

heheh it'd be nice? what for? :)

OpenStudy (jdoe0001):

dress matching?

OpenStudy (darkbluechocobo):

Because all you would have to say is Dependent and Infinite many solutions

OpenStudy (jdoe0001):

so ok... so now we know they're no "dependent" with infinite solutions..... so.... what do you think? do they have a solution? have you tried solving it?

OpenStudy (darkbluechocobo):

How O_O i don't know how D:

OpenStudy (darkbluechocobo):

*flips table*

OpenStudy (jdoe0001):

well. notice the 1st equation, it's saying that y = x+4 so if y = x + 4 let us use THAT in the 2nd equation 5y = 5x + 35 but we know that y = x+4 thus 5( x + 4 ) = 5x + 35 # now solve for "x"

OpenStudy (darkbluechocobo):

they do not equal

OpenStudy (darkbluechocobo):

o wait

OpenStudy (jdoe0001):

hmmm what do you get?

OpenStudy (darkbluechocobo):

so 5x-5x is 0x

OpenStudy (darkbluechocobo):

x=15?

OpenStudy (darkbluechocobo):

now I am assuming you fill in x to solve for y

OpenStudy (darkbluechocobo):

y=19 x=15

OpenStudy (darkbluechocobo):

i think i screwed up

OpenStudy (jdoe0001):

ahhh.... I see , so there's NO SOLUTION lemme make it simpler lemme solve for "y" on the 2nd equation \(\bf 5y = 5x + 35 \implies y = \cfrac{5x}{5}+\cfrac{35}{5}\implies y = x + 7\) so notice, their slope, same slope, \(\bf y = \color{red}{1}x + 4 \qquad y = \color{red}{1}x+7\) what does that mean? parallel lines have the same slope, thus they never meet a solution to a system of equations is when the 2 equations meet, if they don't, well, there's no solution thus the system is INCONSISTENT

OpenStudy (darkbluechocobo):

o.o Oooooooo I see

OpenStudy (darkbluechocobo):

Thank you for explaining it o:

OpenStudy (darkbluechocobo):

Do you think you could help me with more kind person o.o

OpenStudy (darkbluechocobo):

Thank you also may I add the Tunnel has light in my mind

OpenStudy (darkbluechocobo):

o:

OpenStudy (darkbluechocobo):

Thank you that is the website of my dreams

OpenStudy (darkbluechocobo):

Classify the system and determine the number of solutions. Check all that apply. 7x + y = 13 28x + 4y = -12 consistent infinitely many solutions inconsistent no solution So with this one would you do the same thing?

OpenStudy (darkbluechocobo):

so y=-7x+13

OpenStudy (jdoe0001):

do the same, solve for "y" on each case to check their slope

OpenStudy (darkbluechocobo):

28x + 4( -7x + 13)=-12

OpenStudy (jdoe0001):

well. you don't need to solve it, just to check if it has a solution

OpenStudy (darkbluechocobo):

28x +-28x + 52= -12

OpenStudy (darkbluechocobo):

Oh

OpenStudy (darkbluechocobo):

x=-64

OpenStudy (jdoe0001):

solving the 2nd equation for "y", gives us y = -7x -3

OpenStudy (jdoe0001):

notice their slopes \(\bf y=\color{red}{-7}x+13 \qquad y = \color{red}{-7}x -3\)

OpenStudy (darkbluechocobo):

oooo -7x is the slope and -3 is the center point

OpenStudy (darkbluechocobo):

same slope

OpenStudy (jdoe0001):

yeap, that means the graph is just 2 parallel lines, they never meet, thus no solution, thus inconsistent

OpenStudy (darkbluechocobo):

Is this inconsistent again then o.o

OpenStudy (darkbluechocobo):

You beat me too it :P

OpenStudy (jdoe0001):

har har

OpenStudy (darkbluechocobo):

So it is just inconsistent right?

OpenStudy (jdoe0001):

yeap

OpenStudy (darkbluechocobo):

Thank you again :3

OpenStudy (jdoe0001):

yw

OpenStudy (darkbluechocobo):

Classify the system and determine the number of solutions.Check all that apply. $8y - 24x = 64$ $9y + 45x = 72$ inconsistent infinitely many solutions consistent dependent one solution independent

OpenStudy (darkbluechocobo):

there is one more after this one